7. Show that V = R2 with the given operations and o is not a vector space. 11 [][i:1-1 3]..[]-[] Y1
Answer #1

Consider V=\mathbb{R}^2 With the operations \oplus\; and \; \odot defined as.

\begin{bmatrix} x_1\\ y_1 \end{bmatrix} \oplus \begin{bmatrix} x_2\\ y_2 \end{bmatrix}=\begin{bmatrix} x_1-x_2\\ y_1 -y_2 \end{bmatrix}, c\; \odot \begin{bmatrix} x_1\\ y_1 \end{bmatrix}=\begin{bmatrix} cx_1\\ cy_1 \end{bmatrix}

Consider two vectors

u=\begin{bmatrix} 0\\ 0 \end{bmatrix} \; and \; v= \begin{bmatrix} 1\\ 1 \end{bmatrix} \; in \; \mathbb{R}^2

Then we have

u\oplus v=\begin{bmatrix} 0\\ 0 \end{bmatrix}\oplus \begin{bmatrix} 1\\ 1 \end{bmatrix}=\begin{bmatrix} 0-1\\ 0-1 \end{bmatrix}=\begin{bmatrix} -1\\ -1 \end{bmatrix}

and

v\oplus u=\begin{bmatrix} 1\\ 1 \end{bmatrix}\oplus \begin{bmatrix} 0\\ 0 \end{bmatrix}=\begin{bmatrix} 1-0\\ 1-0 \end{bmatrix}=\begin{bmatrix} 1\\ 1 \end{bmatrix}

It is observed that

u\oplus v\ne v\oplus u

Therefore  \oplus is not commutative.

Hence V is not a vector space.